OpenStudy (anonymous):

a shell thrown towards a target from the ground with a fixed velocity falls 'p' metres beyond the target when fired at an angel 'a' & falls 'q' meters short when fired at an angle 'b' . show that the correct angle of projection is 1/2xsin^-1 ((p.sin2a +q.sin2b)/(p+q))

OpenStudy (jamesj):

Just to make sure: the correct angle theta satisfies \[\frac{1}{\theta} = 2.\arcsin \left( \frac{p.\sin 2a + q.\sin 2b}{p + q} \right)\] ? Nice problem.

OpenStudy (jamesj):

No ... it is that theta equals 1/2 times the arcsin expression

OpenStudy (anonymous):

no 2x\[\theta =\]sin^-1 ((p.sin2a +q.sin2b)/(p+q))

OpenStudy (jamesj):

yes. This actually isn't so bad. Suppose a projectile is fired with velocity v at an angle alpha to the horizontal. What is the expression for how far the projectile travels along the horizontal?

OpenStudy (anonymous):

\[R= u ^{2}\sin2\alpha/ g\]

OpenStudy (jamesj):

The horizontal velocity is v.cos(alpha) and hence it travels v.cos(alpha).T where T is the time the projectile reaches the horizontal line again. We can find T from considering the vertical velocity, call that vy(t) vy(t) = v.sin(alpha) - gt hence vertical distance, dy(t) by integrating and setting dy(0) = 0 dy(t) = v.sin(alpha)t - gt^2/2 We now find T, as it satisfies dy(T) = 0. So do that and find the expression for the distance travelled by the projectile, call it d(alpha)

OpenStudy (jamesj):

Right.

OpenStudy (jamesj):

Now you have d(a) = x - p and d(b) = x + q where x is the distance we want. Now manipulate these equations and find the theta such that d(theta) = x.